# Hole spin qubits in unstrained Germanium layers

### v1

@@ I Introduction
It would, therefore, be desirable to increase the HH/LH mixing and reduce the $g$-factor anisotropy. As discussed in Ref. [32], this may be achieved with strain engineering, but a scalable design is still lacking. An alternative solution is to host the qubits in a bulk Ge substrate insulated from the gate stack by a thin, strained GeSi barrier [33, 34]. The quantum dots are then accumulated at the Ge/GeSi interface by the electric field from the gates. As the Ge substrate is unstrained, the HH/LH band gap is expected significantly smaller and the HH/LH mixing much stronger.
In this work, we explore the prospects for such unstrained bulk Ge qubits with detailed numerical simulations. We analyze the dependence of the $g$-factors on the electrical confinement, and show that the $g$-factor anisotropy can indeed be significantly reduced even for moderate HH/LH mixings. The stronger mixing increases the Rabi frequencies $f_{\mathrm{R}}$ but decreases the dephasing time $T_{2}^{*}$; nonetheless, the quality factor $Q_{2}^{*}=2f_{\mathrm{R}}T_{2}^{*}$ is larger than in strained heterostructures. Most importantly, the dependence of these quantities on the magnetic field orientation is much broadened, allowing for an easier optimization of the operating point in many-qubit systems. We discuss the implications for the development of hole spin qubit technologies.
@@ II.2 Dimensions and $g$-factors of the dots
As expected, the vertical extension of the dot increases with $L_{w}$, but is ultimately limited by the vertical electric field from the gates that tends to squeeze the hole at the top Ge/Ge0.8Si0.2 interface. Indeed, $\ell_{\perp}\approx 0.18L_{w}$ is consistent with a square well model for small $L_{w}$, but departs from this trend in thick Ge films. There is little difference between strained and unstrained wells as the biaxial strain has almost no effect on the vertical confinement mass. The dot also extends laterally with increasing $L_{w}$, primarily because the in-plane (transport) mass of the hole decreases (from $m_{\parallel}\approx 0.08\,m_{0}$ for $L_{w}=10$ nm to $m_{\parallel}\approx 0.06\,m_{0}$ for $L_{w}=70$ nm, with $m_{0}$ the free electron mass)  [^2] To lowest order, the transport mass is $m_{\parallel}\approx m_{0}/(\gamma_{1}+\gamma_{2}-\gamma_{h})$, where $\gamma_{1}=13.18$, $\gamma_{2}=4.24$ and $\gamma_{h}$ is defined after Eq. (1) [42].. This is also why the dots are more localized in the unstrained Ge wells that show heavier transport masses ($m_{\parallel}\approx 0.11\,m_{0}$ for $L_{w}=10$ nm and $m_{\parallel}\approx 0.08\,m_{0}$ for $L_{w}=70$ nm). Although undesirable, the enhanced localization in unstrained Ge wells remains limited.
@@ II.3 Spin manipulation
This matrix is dominated by the $g^{\prime}_{xz}$ and $g^{\prime}_{zx}$ terms that capture the rotations of the principal axes of the $g$-matrix in the inhomogeneous electric field of the L gate [35] (and also the effects of a weaker cubic Rashba interaction [47, 48]). Namely, the axis of $z^{\prime}$ of the $g_{\perp}$-factor (and the orthogonal, equatorial $(x^{\prime}y^{\prime})$ plane) rock from left to right when the dot is driven by the L gate, which tilts the precession axis of the spin and results in coherent spin rotations at resonance. This mechanism gives rise to the prominent peak for in-plane magnetic fields, because the effects of small excursions of $\mathbf{B}$ around the effective equatorial $(x^{\prime}y^{\prime})$ plane are amplified by $g_{\perp}>g_{\parallel}$, and because the Rabi frequency is $\propto B$, which is larger in-plane at constant Larmor frequency.
@@ III Discussion
NOTE
In the final stage of submitting this manuscript a preprint demonstrating a low-disorder two-dimensional hole gas in an unstrained Ge/strained SiGe heterojunction has appeared [51].
@@ ACKNOWLEDGEMENTS
This work was supported by the “France 2030” program (PEPR PRESQUILE-ANR-22-PETQ-0002).
DATA AVAILABILITY
Data is available from the corresponding author upon reasonable request.
AUTHORS CONTRIBUTIONS
L. M., M. J. R. and E. A. R. M. performed the calculations. Y. M. N. initiated and supervised the work. All authors contributed to the analysis of the results and review of the manuscript.
COMPETING INTERESTS
The authors declare no competing interests.

### v2

@@ I Introduction
It would, therefore, be desirable to increase the HH/LH mixing and reduce the $g$-factor anisotropy. As discussed in Ref. [32], this may be achieved with strain engineering, but a scalable design is still lacking. An alternative solution is to host the qubits in a bulk Ge substrate insulated from the gate stack by a thin, strained GeSi barrier [33, 34]. The quantum dots are then accumulated at the Ge/GeSi interface by the electric field from the gates. As the Ge substrate is unstrained, the HH/LH band gap is expected significantly smaller and the HH/LH mixing much stronger [35]. The growth of such a structure, with the formation of a high mobility hole gas at the interface, has actually been demonstrated very recently [36].
In this work, we explore the prospects for unstrained bulk Ge qubits with detailed numerical simulations. We analyze the dependence of the $g$-factors on the electrical confinement, and show that the $g$-factor anisotropy can indeed be significantly reduced even for moderate HH/LH mixings. The stronger mixing increases the Rabi frequencies $f_{\mathrm{R}}$ but decreases the dephasing time $T_{2}^{*}$; nonetheless, the quality factor $Q_{2}^{*}=2f_{\mathrm{R}}T_{2}^{*}$ is larger than in strained heterostructures. Most importantly, the dependence of these quantities on the magnetic field orientation is much broadened, allowing for an easier optimization of the operating point in many-qubit systems. We discuss the implications for the development of hole spin qubit technologies.
@@ II.2 Dimensions and $g$-factors of the dots
As expected, the vertical extension of the dot increases with $L_{w}$, but is ultimately limited by the vertical electric field from the gates that tends to squeeze the hole at the top Ge/Ge0.8Si0.2 interface. Indeed, $\ell_{\perp}\approx 0.18L_{w}$ is consistent with a square well model for small $L_{w}$, but departs from this trend in thick Ge films. There is little difference between strained and unstrained wells as the biaxial strain has almost no effect on the vertical confinement mass. The dot also extends laterally with increasing $L_{w}$, primarily because the in-plane (transport) mass of the hole decreases (from $m_{\parallel}\approx 0.08\,m_{0}$ for $L_{w}=10$ nm to $m_{\parallel}\approx 0.06\,m_{0}$ for $L_{w}=70$ nm, with $m_{0}$ the free electron mass)  [^2] To lowest order, the transport mass is $m_{\parallel}\approx m_{0}/(\gamma_{1}+\gamma_{2}-\gamma_{h})$, where $\gamma_{1}=13.18$, $\gamma_{2}=4.24$ and $\gamma_{h}$ is defined after Eq. (1) [44].. This is also why the dots are more localized in the unstrained Ge wells that show heavier transport masses ($m_{\parallel}\approx 0.1\,m_{0}$ for $L_{w}=10$ nm and $m_{\parallel}\approx 0.08\,m_{0}$ for $L_{w}=70$ nm). Although undesirable, the enhanced localization in unstrained Ge wells remains limited.
@@ II.3 Spin manipulation
This matrix is dominated by the $g^{\prime}_{xz}$ and $g^{\prime}_{zx}$ terms that capture the rotations of the principal axes of the $g$-matrix in the inhomogeneous electric field of the L gate [37] (and also the effects of a cubic Rashba interaction [49, 50, 35]). Namely, the axis of $z^{\prime}$ of the $g_{\perp}$-factor (and the orthogonal, equatorial $(x^{\prime}y^{\prime})$ plane) rock from left to right when the dot is driven by the L gate, which tilts the precession axis of the spin and results in coherent spin rotations at resonance. This mechanism gives rise to the prominent peak for in-plane magnetic fields, because the effects of small excursions of $\mathbf{B}$ around the effective equatorial $(x^{\prime}y^{\prime})$ plane are amplified by $g_{\perp}>g_{\parallel}$, and because the Rabi frequency is $\propto B$, which is larger in-plane at constant Larmor frequency.
@@ ACKNOWLEDGEMENTS
This work was supported by the “France 2030” program (PEPR PRESQUILE-ANR-22-PETQ-0002) and by the Horizon Europe Framework Program (grant agreement 101174557 QLSI2).